OpenStudy (anonymous):

find the binomial expansion of (x-1/x^3)^10

OpenStudy (camper4834):

oh well lets start with the first rule about exponents

OpenStudy (anonymous):

x^10-10x^6+45x^2-120/x^2+210/x^6-252/x^10 +210/x^14 -120/x^18+45/x^22-10/x^26+1/x^30

OpenStudy (anonymous):

okay i seen the answer too in wolfram alpha i need to understand not the answer lol

OpenStudy (anonymous):

Binomial theorem: \[\large(a+b)^n=\sum_{k=0}^na^kb^{n-k}\] Set \(a=x\) and \(b=-\dfrac{1}{x^3}\).

OpenStudy (anonymous):

Oh and \(n=10\).

OpenStudy (anonymous):

okay so if its x-1/x^3

OpenStudy (anonymous):

i should jus rearrange it to be -1/x^3 + x correct

OpenStudy (anonymous):

That shouldn't matter.

OpenStudy (anonymous):

\((a+b)^n=(b+a)^n\)

OpenStudy (anonymous):

hmmm ill write that down ..

OpenStudy (anonymous):

how will i get the coefficents though

OpenStudy (anonymous):

That's where the binomial coefficient comes in. The first term, for \(k=0\), would be \[\binom{10}0x^0\left(-\frac{1}{x^3}\right)^{10-0}=\frac{10!}{0!(10-0)!}\times1\times\left(-\frac{1}{x^3}\right)^{10}=\frac{1}{x^{30}}\]

OpenStudy (anonymous):

okay thx so the coefficient is just (n-1 k ) right

OpenStudy (anonymous):

oh no its

OpenStudy (anonymous):

Not sure where you're getting \(n-1\) from. Here's the def. of the binomial coefficient: \[\binom nk=\frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

yea i saw it was wrong was thinking something else

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw